OpenStudy (anonymous):

how do I re-write the following expression m as the product of three terms: m=e^(2x+3y-4z)

OpenStudy (lucaz):

you can use the property \[\huge{a^{(x+y+z)}=a^x\times a^y\times a^z}\]

OpenStudy (lucaz):

when you have the same base in a multiplication you can keep the base and add the exponents

OpenStudy (anonymous):

Thank you @lucaz. My teacher was saying something about the inverse function of ln with this answer. She said there's more to it. Do you know what she's talking about?

OpenStudy (lucaz):

well, log is the inverse function of the exponential function

OpenStudy (lucaz):

natural log has e as it's base

OpenStudy (anonymous):

@lucaz okay. So, place ln in front of each exponential function?

OpenStudy (lucaz):

@texaschic101 can you help here?

OpenStudy (lucaz):

let's see if someone else could explain it better

OpenStudy (anonymous):

okay thank you

OpenStudy (lucaz):